Rectangle closet: If the perimeter of a rectangular sign is 44cm and the width is 2cm shorter than half the length, then what is the length and width?
Answer by mananth(16946)

let length be x cm
width = (x/2)-2 cm
perimeter = 2*(L+(x/2))
Perimeter = 44 cm
44=2*((x)+(x/2 -2))
44= 2(3x/2 -2 )
/2
22= 3x/2 -2+2
24=3x/2
multiply by 2
48=3x
/3
16cm =x the length
width = 16/2 - 2
=6 cm
....